# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ Abstract
Integrated sensing and communication (ISAC) is envisioned as a key technology for future sixth-generation (6G) networks. Classical ISAC system considering monostatic and/or bistatic settings will inevitably degrade both communication and sensing performance due to the limited service coverage and easily blocked transmission paths. Besides, existing ISAC studies usually focus on downlink (DL) or uplink (UL) communication demands and unable to achieve the systematic DL and UL communication tasks. These challenges can be overcome by networked FD ISAC framework. Moreover, ISAC generally considers the trade-off between communication and sensing, unavoidably leading to a loss in communication performance. This shortcoming can be solved by the emerging movable antenna (MA) technology. In this paper, we utilize the MA to promote communication capability with guaranteed sensing performance via jointly designing beamforming, power allocation, receiving filters and MA configuration towards maximizing sum rate. The optimization problem is highly difficult due to the unique channel model deriving from the MA. To resolve this challenge, via leveraging the cutting-the-edge majorization-minimization (MM) method, we develop an efficient solution that optimizes all variables via convex optimization techniques. Extensive simulation results verify the effectiveness of our proposed algorithms and demonstrate the substantial performance promotion by deploying MA in the networked FD ISAC system.
Index Terms:
networked integrated sensing and communication (ISAC), full-duplex (FD), movable antenna (MA), majorization-minimization (MM) algorithm.
@@ I Introduction
Recently, with the rapid increase in devices requiring precision sensing and efficient communication, integrated sensing and communication (ISAC) [1]--[2] technologies have attracted great interest from both industry and academia. In this context, ISAC is envisioned as a promising solution, which aiming to deploy both communication and sensing functionalities on one unified hardware platform which allows for the sharing of frequency spectrums, hardware facilities, and signal-processing units. Many recent advancements in the joint design of radar sensing and communication are extensively documented in [1]--[2] and their associated references.
However, the conventional fixed-position antennas (FPAs) at the base station (BS) and/or mobile user in the ISAC system generally fail to fully exploit the degrees of freedom (DoFs) in the continuous spatial domain, leading to performance loss in both sensing and communication tasks. A novel movable antenna (MA) technology [3]--[4] is expected to overcome the this limitation. The MA is considered as a viable approach for enhancing wireless network performance, linked to a radio frequency (RF) chain through a flexible cable. It can reconstruct channel conditions by flexibly adjusting its position in a given spatial area with the aid of a driver component or by other means. The various aspects of wireless systems with applying MA have been extensively studied recently, see [5]--[9] and their references.
@@ I-B Motivations and Contributions
As shown previously, although extensive literature has explored the joint beamforming optimization for MA-aided ISAC systems, most studies have focused on monostatic and/or bistatic settings. These settings have limited service coverage and insufficient both sensing and communication performance in complex environments. In contrast, the networked ISAC is a promising solution to overcome the above shortcomings [1]--[2]. Moreover, the recent literature [10]--[18] only considers the DL communication in MA-aided ISAC networks. Note the waveform design in networked FD ISAC system remains largely unexplored, apart from the recent study [19]. More importantly, all the existing works [10]--[19] have not taken into account the integration of MA into the networked FD ISAC system. Therefore, we are motivated to study an MA-aided networked FD ISAC system to fully enhance its communication and sensing capabilities. Specifically, the contributions of this paper are elaborated as follows:
@@
  • This paper considers the joint beamforming and position design in a networked FD ISAC system enabled by MA to promote simultaneous DL and uplink (UL) communication and target sensing. This novel framework can i) effectively enlarge the line-of-sight (LoS) region for sensing; ii) provide concurrent DL and UL communication transmission; iii) reconstruct the wireless channel condition by the MA technology.
@@
  • We investigate the maximization of the sum-rate for both DL and UL users while guaranteeing radar sensing quality via jointly optimizing the BSs’ beamforming, UL users power allocation, receive filter and MAs’ positions. To the best of our knowledge, this problem has not been considered in the existing literature, e.g., [10]--[19].
@@
  • The joint beamforming and MA configuration design to improve sum-rate of both DL and UL users is a highly non-convex problem due to the position coefficients of MA. Based on the cutting-the-edge majorization-minimization (MM) transformation [20], we successfully develop an effective algorithm that updates the position coefficients of MA by convex optimization techniques.
@@
  • Extensive numerical results are provided to verify the effectiveness of our proposed solution. Meanwhile, experiment results validate that the deployment of MA can significantly benefit both DL and UL communication, promoting the sum-rate by up to 20.0% compared to the FPA benchmark, in the networked FD ISAC scheme.
@@ II-A System Model
Figure 1: A MA-aided networked FD ISAC system. (a): System model (b): Illustrations of the MAs at the BS and users, respectively.
@@
As shown in Fig. 1, we consider an MA-enhanced networked FD ISAC system that comprises $M_{t}$ transmit BSs (TBSs), $M_{r}$ receive BSs (RBSs), $K_{d}$ single MA DL mobile users, $K_{u}$ single MA UL mobile users, one point-like sensing target, and $K_{t}$ clutter sources. Moreover, TBSs and RBSs are respectively equipped with $N_{t}$ and $N_{r}$ MAs. Let $\mathcal{M}_{t}\triangleq\{1,\cdots,M_{t}\}$, $\mathcal{M}_{r}\triangleq\{1,\cdots,M_{r}\}$, $\mathcal{N}_{t}\triangleq\{1,\cdots,N_{t}\}$, $\mathcal{N}_{r}\triangleq\{1,\cdots,N_{r}\}$, $\mathcal{K}_{d}\triangleq\{1,\cdots,K_{d}\}$, $\mathcal{K}_{u}\triangleq\{1,\cdots,K_{u}\}$ and $\mathcal{K}_{t}\triangleq\{1,\cdots,K_{t}\}$ denoted the sets of TBSs, RBSs, TBSs’ MAs, RBSs’ MAs, DL users, UL users and clutter sources, respectively.
In this network, TBSs, with aid of MA, service DL users and send probing waveform to detect the target, and RBSs, with the assistance of MA, receive the information from UL users and echo signal from both the target and clutters. Besides, we assumed that all TBSs and RBSs achieve perfect synchronization and joint signal processing facilitated by the central controller [21].
@@
Moreover, the MAs are connected to RF chains by flexible cables, and can move within local region in real time [3], [4]. The position of the $n$-th MA of $m$-th BS is given by $\mathbf{t}^{\iota}_{m,n}=[x^{\iota}_{m,n},y^{\iota}_{m,n}]^{T}\in\mathcal{C}$, where $\iota\in\{0,1\}$, index 00 represents $\forall n\in\mathcal{N}_{t}$ and $\forall m\in\mathcal{M}_{t}$, and index $1$ denotes $\forall n\in\mathcal{N}_{r}$ and $\forall m\in\mathcal{M}_{r}$. The positions of single MA at the $k$-th DL user and $l$-th UL user are respectively written as $\mathbf{r}_{0,k}=[x^{d}_{k},y^{d}_{k}]^{T}\in\mathcal{C}$ and $\mathbf{r}_{1,l}=[x^{u}_{l},y^{u}_{l}]^{T}\in\mathcal{C}$, where $\mathcal{C}$ denotes the given two-dimensional (2D) moving square region with a size of $A\times A$ for both the BSs’ and users’ MAs.
@@
Since the MA’s moving area is usually much smaller than the signal propagation distance, the links between the BSs/users and the users are assumed to follow the far-field wireless channel model [3], [4]. Therefore, changing the positions of MAs has no effect on the angle of departure (AoD), the angle of arrival (AoA), and the amplitude of the complex path coefficient, while only the phase of the complex path coefficient varies with changes in MAs’ position.
@@
Let $L^{t,\iota}_{m,k}$ and $L^{r,\iota}_{m,k}$ denote the total number of transmit and receive channel paths at the $m$-th BS from $k$-th user, respectively. And then, $L^{t,2}_{k,l}$ and $L^{r,2}_{k,l}$ are respectively denoted as the total number of transmit and receive channel paths between DL user $k$ and UL user $l$. The elevation and azimuth AoDs of $i$-th transmit path between BS $m$ and user $k$ are respectively denoted as $\theta^{t,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{t,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$. Similarly, the elevation and azimuth AoAs of $i$-th receive path between BS $m$ and user $k$ are represented as $\theta^{r,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{r,\iota}_{m,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, respectively. Besides, denote the elevation and azimuth AoDs of $i$-th transmit path between UL user $l$ and DL user $k$ as $\theta^{t,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{t,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, respectively, and denote separately the elevation and azimuth AoAs of $i$-th receive path between UL user $l$ and DL user $k$ as $\theta^{r,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi^{r,2}_{l,k,i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$. Next, the difference in signal propagation for the $i$-th transmit path of $k$-th user between the position of $n$-th MA of $m$-th BS and its reference point, i.e., $\mathbf{t}^{\iota}_{m,0}=[0,0]^{T}$, can be given as
$\rho^{t,\iota}_{m,k,n,i}(\mathbf{t}^{\iota}_{m,n})=({\mathbf{t}_{m,n}^{{\iota}}})^{T}\mathbf{a}^{t,\iota}_{m,k,i}$ (1)
$=x^{\iota}_{m,n}\cos(\theta^{t,\iota}_{m,k,i})\sin(\phi^{t,\iota}_{m,k,i})+y^{\iota}_{m,n}\sin(\theta^{t,\iota}_{m,k,i}),$
@@
and its phase difference can be obtained as $2\pi\rho^{t,\iota}_{m,k,n,i}(\mathbf{t}^{\iota}_{m,n})/\lambda$, where $\lambda$ denotes the carrier wavelength. Thus, the field-response vector (FRV) of the transmit channel paths between $k$-th user and $m$-th BS’ $n$-th MA is given by
$\mathbf{h}^{\iota}_{m,k,n}\!\!=\!\![e^{j\frac{2\pi}{\lambda}\rho^{t,\iota}_{m,k,n,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,\iota}_{m,k,n,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,\iota}_{m,k,n,L^{t,\iota}_{m,k}}}]^{T},$ (2)
and then the field response matrixs (FRMs) of all $N_{t}$ MAs for $m$-th TBS and all $N_{r}$ MAs for $p$-th TBS can be written as
@@
$\mathbf{H}^{1}_{p,l}({\mathbf{t}^{1}_{p}})=[\mathbf{h}^{1}_{p,l,1},\mathbf{h}^{1}_{p,l,2},\cdots,\mathbf{h}^{1}_{p,l,N_{r}}]\in\mathbb{C}^{L^{t,1}_{p,l}\times N_{r}},$ (4)
@@
respectively, where $\mathbf{t}^{0}_{m}\triangleq\{\mathbf{t}^{0}_{m,n}\}$ and $\mathbf{t}^{1}_{p}\triangleq\{\mathbf{t}^{1}_{p,i}\}$. Moreover, the FRVs of the receive channel paths between $k$-th DL user and $m$-th TBS and between $l$-th UL user and $p$-th RBS are respectively defined as
$\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,L^{r,0}_{m,k}}}],$ (5)
$\mathbf{h}^{0,1}_{p,l}(\mathbf{r}_{1,l})=[e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{p,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{p,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{p,l,L^{r,1}_{p,l}}}],$ (6)
where $\rho^{r,0}_{m,k,i}(\mathbf{r}_{0,k})\!\triangleq\!\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,0}_{m,k,i}\!=\!x^{d}_{k}\cos(\theta^{r,0}_{m,k,i})\sin(\phi^{r,0}_{m,k,i})+y^{d}_{k}\sin(\theta^{r,0}_{m,k,i})$ and $\rho^{r,1}_{p,l,j}(\mathbf{r}_{1,l})\!\!\triangleq\!\!\mathbf{r}_{1,l}^{T}\mathbf{a}^{r,1}_{p,l,j}\!=\!y^{u}_{l}\sin(\theta^{r,1}_{p,l,j})+x^{u}_{l}\cos(\theta^{r,1}_{p,l,j})\sin(\phi^{r,1}_{p,l,j})$.
Besides, the FRVs of the transmit and receive channel paths between $k$-th DL user and $l$-th UL user are respectively expressed as
$\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l})=[e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,L^{t,2}_{k,l}}}],$ (7)
$\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,L^{r,2}_{k,l}}}],$ (8)
@@
where $\rho^{t,2}_{k,l,i}(\mathbf{r}_{1,l})\triangleq\mathbf{r}_{1,l}^{T}\mathbf{a}^{t,2}_{k,l,i}=x^{u}_{l}\cos(\theta^{t,2}_{k,l,i})\sin(\phi^{t,2}_{k,l,i})+y^{u}_{l}\sin(\theta^{t,2}_{k,l,i})$ and $\rho^{r,2}_{k,l,i}(\mathbf{r}_{0,k})\triangleq\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,2}_{k,l,i}=y^{d}_{k}\sin(\theta^{r,2}_{k,l,i})+x^{d}_{k}\cos(\theta^{r,2}_{k,l,i})\sin(\phi^{r,2}_{k,l,i})$. Furthermore, the matrixes $\mathbf{\Sigma}^{0}_{m,k}$, $\mathbf{\Sigma}^{1}_{p,l}$, and $\mathbf{\Sigma}^{2}_{k,l}$ are respectively denoted as the responses of all transmit and receive paths from the $m$-th TBS to the $k$-th DL user, from the $p$-th RBS to the $l$-th DL user and from the $l$-th UL user to the $k$-th DL user. Therefore, the channel vectors between the $m$-th TBS and the $k$-th DL user and between the $p$-th RBS and the $l$-th UL user are given by
$\mathbf{h}_{d,m,k}(\mathbf{t}^{0}_{m},\mathbf{r}_{0,k})^{H}=\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})^{H}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}(\mathbf{t}^{0}_{m})\in\mathbb{C}^{1\times N_{t}},$ (9)
$\mathbf{h}_{u,p,l}(\mathbf{t}^{1}_{p},\mathbf{r}_{0,l})^{H}=\mathbf{h}^{0,1}_{p,l}(\mathbf{r}_{0,l})^{H}\mathbf{\Sigma}^{1}_{p,l}\mathbf{H}^{1}_{p,l}(\mathbf{t}^{1}_{p})\in\mathbb{C}^{1\times N_{r}},$ (10)

### v3

# Movable Antenna Enhanced Networked Integrated Sensing and Communication System

@@ Abstract
Integrated sensing and communication (ISAC) is a key technology for future 6G networks. Most existing studies focus on monostatic and/or bistatic setups with limited coverage and capabilities. Networked ISAC systems with distributed base stations (BSs) can overcome these limitations. Moreover, movable antenna (MA) architectures offer improved ISAC performance over fixed-position antennas (FPAs) by enabling adaptable antenna movement. In this paper, we utilize the MA to promote communication capability with guaranteed sensing performance via jointly designing beamforming, power allocation, receiving filters and position configuration of transmit/receive MA towards maximizing the sum rate for both downlink (DL) and uplink (UL) users. The optimization problem is highly difficult due to the unique channel model derived from the position coefficient of the MA. To resolve this challenge, via leveraging the cutting-the-edge majorization-minimization (MM) method, we develop an efficient solution that optimizes all variables via convex optimization techniques. Extensive simulation results verify the effectiveness of our proposed algorithms and demonstrate the substantial performance promotion by deploying the MA framework in the networked ISAC system.
@@ I Introduction
Recently, with the rapid increase in devices requiring precision sensing and efficient communication, integrated sensing and communication (ISAC) [1]-[2] technologies have attracted great interest from both industry and academia. In this context, ISAC is envisioned as a promising solution, which aims to deploy both communication and sensing functionalities on one unified hardware platform, allowing the sharing of frequency spectrum, hardware resources, and signal-processing units. Many recent advancements in the joint design of radar sensing and communication are extensively documented in [1]-[2] and their associated references.
However, the conventional fixed-position antenna (FPA) at the base station (BS) and/or mobile user in the ISAC system generally fail to fully exploit the degrees of freedom (DoFs) in the continuous spatial domain, leading to performance loss in both sensing and communication tasks. A novel movable antenna (MA) technology [3]-[5] is expected to overcome the limitation. The MA is considered as a viable approach for enhancing wireless network performance, linked to a radio frequency (RF) chain through a flexible cable. It can reconstruct channel conditions by flexibly adjusting its position in a given spatial area with the aid of a driver component or by other means. The various aspects of wireless systems by applying the MA framework have been extensively studied recently, see [6]-[11] and their references.
@@ I-A Related Works
In the mono-static ISAC system, the full-duplex (FD) operation presents a significant challenge due to self-interference [1], which cannot be overlooked. The work [12] aimed to deal with the challenges associated with the FD operation in mono-static ISAC systems operating at millimeter-Wave (mm-Wave) frequencies, and proposed the beamforming solutions designed to maximize the beamformer power in the sensing direction while simultaneously constraining the beamformer power in the communication directions. The authors of [13] and [14] adopted the Orthogonal Frequency Division Multiplexing (OFDM) waveform to estimate the Direction of Arrival (DoA), range, and relative velocity of the radar targets while maximizing the downlink (DL) communication rate in the mono-static FD ISAC network operating at the mm-Wave frequencies. The paper [15] investigated a mono-static FD ISAC system, in which the BS simultaneously conducts target detection while communicating with multiple DL and uplink (UL) users, sharing the same time-frequency resources. The authors of [16] utilized the FD in the ISAC transceiver to solve the echo-miss problem.
The networked ISAC system with multiple transmitters and/or receivers can significantly improve both the communication and radar sensing performance [1]. In [17], the work investigated a networked ISAC system that considers two types of users: those capable of canceling interference from dedicated sensing signals and those that are not. The paper [18] presented a cell-free ISAC MIMO system, deriving sensing signal-to-noise (SNR) and developing joint beamforming and power allocation for optimal communication and sensing performance. The authors of [19] simultaneously considered the active and passive sensing tasks in a downlink distributed ISAC system, adopting different fusion strategies based on the backhaul capacity and enhancing sensing performance while meeting minimum signal-to-interference-plus-noise (SINR) requirements. The work [20] demonstrated a multi-static ISAC design using network-assisted full-duplex (NAFD) cell-free networks, proposing a deep Q-network for optimizing access point (AP) duplex modes.
@@ I-B Motivations and Contributions
As shown previously, although several studies [12]-[20] have investigated joint beamforming optimization in the ISAC systems, most have relied on traditional FPAs, limiting their adaptability and performance. Since the MA technology can effectively utilize spatial DoFs in practical networks [4]-[5], extensive literature [21]-[33] has explored the joint beamforming and MA’s position optimization for MA-aided ISAC systems but primarily focused on monostatic and/or bistatic configurations. These settings have limited service coverage and inadequate performance in both sensing and communication within complex environments. However, the networked ISAC is a promising solution for enhancing coverage, spectral efficiency, and reliability [1]-[2]. Moreover, these works [21]-[29], [33], primarily concentrate on DL communication, overlooking the crucial aspect of UL communication which is essential for comprehensive ISAC functionality. This oversight can restrict the practical application of existing systems in dynamic environments where bidirectional communication is vital. Therefore, we are motivated to study an MA-aided networked ISAC system, that incorporates multi-static configurations and accommodates both DL and UL users, aiming to enhance its communication and sensing capabilities. Specifically, the contributions of this paper are elaborated as follows:
@@
  • This paper considers the joint beamforming and antennas’ position design in a networked ISAC system enabled by the MA structure to promote simultaneous DL and UL communication and target sensing functionalities, which has not yet been thoroughly investigated. This novel framework can i) effectively enlarge the service region for both communication and sensing; ii) supply more spatial DoFs by leveraging the MA technology to reconstruct the wireless channel condition; iii) provide comprehensive bidirectional communication (i.e., DL and UL communication) transmission.
@@
  • We investigate the maximization of the sum-rate for both DL and UL users while guaranteeing target detection quality via jointly optimizing the BSs’ beamforming, UL users’ power allocation, receive filter and MAs’ positions. Besides, this paper considers a very generic signal propagation model, which fully takes into account both downlink and uplink communication and reflected radar probing signals. As will be seen, this consideration in modeling significantly complicates the beamforming and antennas’ position design task. To the best of our knowledge, this problem has not been considered in the existing literature, e.g., [21]-[33].
@@
  • The joint beamforming and MA position’s configuration design to improve sum-rate of both DL and UL users is a highly non-convex problem due to the position coefficients of MA. Via utilizing cutting-the-edge optimization frameworks including the weighted minimum mean squared error (WMMSE) [34] and majorization-minimization (MM) [35], we have developed a closed solution for updating MA’s position of the DL users. Furthermore, we successfully develop an effective algorithm that updates the position coefficients of MA by convex optimization techniques.
@@
  • Last but not least, extensive numerical results are provided to verify the effectiveness of our proposed solution, which demonstrates as: 1) our proposed algorithm can converge within several iterations; 2) the proposed algorithm can achieve much better than the benchmark cases (especially its counterpart with the FPA-based array at both BSs and users) in terms of sum-rate of both DL and UL users under various settings. Experiment results validate that the deployment of MA can significantly benefit communication performance in the networked ISAC system.
@@ II System Model and Problem Formulation
Figure 1: A MA-aided networked ISAC system. (a): System model (b): Illustrations of the MAs at the BS and users, respectively.
@@
As shown in Fig. 1111The dotted line represents the link between TBS and RBS. , we consider an MA-enhanced networked ISAC system that comprises $M_{t}$ transmit BSs (TBSs), $M_{r}$ receive BSs (RBSs), $K_{d}$ single MA DL mobile users, $K_{u}$ single MA UL mobile users, one point-like sensing target, and $K_{t}$ clutter sources. Moreover, TBSs and RBSs are respectively equipped with $N_{t}$ and $N_{r}$ MAs. Let $\mathcal{M}_{t}\triangleq\{1,\cdots,M_{t}\}$, $\mathcal{M}_{r}\triangleq\{1,\cdots,M_{r}\}$, $\mathcal{N}_{t}\triangleq\{1,\cdots,N_{t}\}$, $\mathcal{N}_{r}\triangleq\{1,\cdots,N_{r}\}$, $\mathcal{K}_{d}\triangleq\{1,\cdots,K_{d}\}$, $\mathcal{K}_{u}\triangleq\{1,\cdots,K_{u}\}$ and $\mathcal{K}_{t}\triangleq\{1,\cdots,K_{t}\}$ denoted the sets of TBSs, RBSs, TBSs’ MAs, RBSs’ MAs, DL users, UL users and clutter sources, respectively.
@@
Moreover, the MAs are connected to RF chains by flexible cables  [^2] The response times of motor-based mechanisms is the order of milliseconds to seconds [3]., and can move within local region in real time [4]-[5]. Let $\mathbf{t}^{0}_{m,n}=[x^{0}_{m,n},y^{0}_{m,n}]^{T}\in\mathcal{C}$ denote the position of the $n$-th MA of $m$-th TBS. Similarly, the $n$-th MA’ position of $m$-th RBS is given by $\mathbf{t}^{1}_{m,n}=[x^{1}_{m,n},y^{1}_{m,n}]^{T}\in\mathcal{C}$. The positions of MA at the $k$-th DL and $l$-th UL users are respectively written as $\mathbf{r}_{0,k}=[x^{d}_{k},y^{d}_{k}]^{T}\in\mathcal{C}$ and $\mathbf{r}_{1,l}=[x^{u}_{l},y^{u}_{l}]^{T}\in\mathcal{C}$, where $\mathcal{C}$ denotes the given two-dimensional (2D) moving square region with a size of $A\times A$ for both the BSs’ and users’ MAs, while the corresponding reference point’s position is $0,0]^{T}$.
In this network, target detection and communication are performed in a time-division manner [1]. TBSs, with the aid of MA, service DL users and send probing waveform to detect the target, and RBSs, with the assistance of MA, receive the information from UL users and the echo signal from the sensing target. We assumed that the BSs are connected to the central controller via fronthaul links and that cooperative sensing of the target through data-level fusion [17]-[18]. All TBSs and RBSs can achieve perfect synchronization and joint signal processing facilitated by the central controller, and the signals transmitted by the TBSs can be made available at all RBSs. Moreover, the link between BSs (i.e., intra-BS channel) is assumed to be quasi-static. Therefore, the received signal from the direct link between the TBS and the RBS can be perfectly eliminated using prior knowledge obtained through the fronthaul links, i.e., passive optical network (PON), coax, cable [36]. Since the users are located in a complex environment with multiple reflected and scattered paths (containing the target/clutters-users links), the channels of the users are assumed to be multiple paths channel model [5], that can be estimated via least square method [3].
II-A Channel Model
@@
Since the MA’s moving area is usually much smaller than the signal propagation distance, the channels between the BSs/users and the users are assumed to follow the far-field wireless channel model [4]-[5]. Therefore, changing the position of MA has no effect on the angle of departure (AoD), the angle of arrival (AoA), and the amplitude of the complex path coefficient, while only the phase of the complex path coefficient varies with changes in MA’s position.
a) Communication Channel:
@@
1) TBS-DL User Channel: Let $L^{t,0}_{m,k}$ and $L^{r,0}_{m,k}$ denote the total number of transmit and receive channel paths at the $m$-th TBS from the $k$-th DL user, respectively. The elevation and azimuth AoDs of the $i$-th transmit path between TBS $m$ and DL user $k$ are respectively denoted as $\theta^{t,0}_{m,k,i}$ and $\phi^{t,0}_{m,k,i}$, while the corresponding elevation and azimuth AoAs of the $i$-th receive path are represented as $\theta^{r,0}_{m,k,i}$ and $\phi^{r,0}_{m,k,i}$, respectively. In the $i$-th transmit path for $m$-th TBS-$k$-th DL user link, the propagation distance difference between the $n$-th MA and its reference point can be given as
$\rho^{t,0}_{m,k,n,i}(\mathbf{t}^{0}_{m,n})=({\mathbf{t}_{m,n}^{{0}}})^{T}\mathbf{a}^{t,0}_{m,k,i}$ (1)
$=x^{0}_{m,n}\cos(\theta^{t,0}_{m,k,i})\sin(\phi^{t,0}_{m,k,i})+y^{0}_{m,n}\sin(\theta^{t,0}_{m,k,i}),$
@@
and its phase difference can be obtained as $2\pi\rho^{t,0}_{m,k,n,i}(\mathbf{t}^{0}_{m,n})/\lambda$, where $\lambda$ denotes the carrier wavelength.
Thus, the field-response vector (FRV) of the transmit channel paths between the $k$-th DL user and the $m$-th TBS’ $n$-th MA is given by
$\mathbf{h}^{0}_{m,k,n}\!\!=\!\![e^{j\frac{2\pi}{\lambda}\rho^{t,0}_{m,k,n,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,0}_{m,k,n,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,0}_{m,k,n,L^{t,0}_{m,k}}}]^{T},$ (2)
and then the corresponding field response matrix (FRM) of all $N_{t}$ MAs can be written as
@@
where $\mathbf{t}^{0}_{m}\triangleq\{\mathbf{t}^{0}_{m,n}\}$. The FRV of the receive channel paths between DL user $k$ and TBS $m$ is defined as
$\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,0}_{m,k,L^{r,0}_{m,k}}}]^{T},$ (4)
@@
where $\rho^{r,0}_{m,k,i}(\mathbf{r}_{0,k})\!\triangleq\!\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,0}_{m,k,i}\!=\!x^{d}_{k}\cos(\theta^{r,0}_{m,k,i})\sin(\phi^{r,0}_{m,k,i})+y^{d}_{k}\sin(\theta^{r,0}_{m,k,i})$. Therefore, the channel vector between the $m$-th TBS and the $k$-th DL user is given by
$\mathbf{h}_{d,m,k}^{H}(\mathbf{t}^{0}_{m},\mathbf{r}_{0,k})=\mathbf{h}^{0,0}_{m,k}(\mathbf{r}_{0,k})^{H}\mathbf{\Sigma}^{0}_{m,k}\mathbf{H}^{0}_{m,k}(\mathbf{t}^{0}_{m})\in\mathbb{C}^{1\times N_{t}},$ (5)
where the matrix $\mathbf{\Sigma}^{0}_{m,k}$ is denoted as the response of all transmit and receive paths from TBS $m$ to DL user $k$.
2) RBS-UL User Channel: Let $L^{t,1}_{m,k}$ and $L^{r,1}_{m,k}$ denote the total number of transmit and receive channel paths at RBS $m$ from UL user $k$, respectively. The elevation and azimuth AoDs of the $i$-th transmit path between RBS $m$ and UL user $k$ are respectively denoted as $\theta^{t,1}_{m,k,i}$ and $\phi^{t,1}_{m,k,i}$, while the corresponding elevation and azimuth AoAs of the $i$-th receive path are represented as $\theta^{r,1}_{m,k,i}$ and $\phi^{r,1}_{m,k,i}$, respectively. In the $i$-th transmit path for $m$-th RBS-$k$-th UL user link, the phase difference of the $n$-th MA and its reference point is given by
$2\pi\rho^{t,1}_{m,k,n,i}(\mathbf{t}^{1}_{m,n})/\lambda=2\pi({\mathbf{t}_{m,n}^{1}})^{T}\mathbf{a}^{t,1}_{m,k,i}/\lambda$ (6)
$=2\pi\big(x^{1}_{m,n}\cos(\theta^{t,1}_{m,k,i})\sin(\phi^{t,1}_{m,k,i})+y^{1}_{m,n}\sin(\theta^{t,1}_{m,k,i})\big)/\lambda.$
Thus, the FRV of the transmit channel paths between the $k$-th UL user and the $m$-th RBS’ $n$-th MA is given by
$\mathbf{h}^{1}_{m,k,n}\!\!=\!\![e^{j\frac{2\pi}{\lambda}\rho^{t,1}_{m,k,n,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,1}_{m,k,n,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,1}_{m,k,n,L^{t,1}_{m,k}}}]^{T},$ (7)
and then the corresponding FRM of all $N_{r}$ MAs is written as
$\mathbf{H}^{1}_{m,k}(\mathbf{t}^{0}_{m})=[\mathbf{h}^{1}_{m,k,1},\mathbf{h}^{1}_{m,k,2},\cdots,\mathbf{h}^{1}_{m,k,N_{r}}]\in\mathbb{C}^{L^{t,1}_{m,k}\times N_{r}},$ (8)
where $\mathbf{t}^{1}_{m}\triangleq\{\mathbf{t}^{1}_{m,n}\}$. The FRV of the receive channel paths between UL user $k$ and RBS $m$ is given as
$\mathbf{h}^{0,1}_{m,k}(\mathbf{r}_{1,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{m,k,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{m,k,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,1}_{m,k,L^{r,1}_{m,k}}}]^{T},$ (9)
where $\rho^{r,1}_{m,k,i}(\mathbf{r}_{1,k})\!\!\triangleq\!\!\mathbf{r}_{1,k}^{T}\mathbf{a}^{r,1}_{m,k,i}\!=\!x^{u}_{k}\cos(\theta^{r,1}_{m,k,i})\sin(\phi^{r,1}_{m,k,i})+y^{u}_{k}\sin(\theta^{r,1}_{m,k,i})$. Therefore, the channel vector between the $m$-th RBS and the $k$-th UL user is represented as
$\mathbf{h}_{u,m,k}^{H}(\mathbf{t}^{1}_{m},\mathbf{r}_{1,k})\!=\!\mathbf{h}^{0,1}_{m,k}(\mathbf{r}_{1,k})^{H}\mathbf{\Sigma}^{1}_{m,k}\mathbf{H}^{1}_{m,k}(\mathbf{t}^{1}_{m})\in\mathbb{C}^{1\times N_{r}},$ (10)
with the matrix $\mathbf{\Sigma}^{1}_{m,k}$ denoting as the response of all transmit and receive paths from the $m$-th RBS to the $k$-th UL user.
3) DL User-UL User Channel: $L^{t,2}_{k,l}$ and $L^{r,2}_{k,l}$ are respectively denoted as the total number of transmit and receive channel paths between DL user $k$ and UL user $l$. Denote the elevation and azimuth AoDs of the $i$-th transmit path between UL user $l$ and DL user $k$ as $\theta^{t,2}_{l,k,i}$ and $\phi^{t,2}_{l,k,i}$, respectively, while the corresponding elevation and azimuth AoAs for the $i$-th receive path are represented as $\theta^{r,2}_{l,k,i}$ and $\phi^{r,2}_{l,k,i}$, respectively.